joan is three times as old as her son richard. ten years from now, joan will be two times as old as richard. how old are they now?
Answer by ikleyn(52776)

.

 x = the son current age;

3x = Joan current age.


In 10 years from now

    3x + 10 = 2*(x+10)

    3x + 10 = 2x + 20

    3x - 2x = 20 - 10

       x    =    10.


ANSWER.  The son is 10 years old.  Joan is 3*10 = 30 years old.

Solved.